# uncompress

### v2

@@ 4. Bounding irrationality of degenerations
In this section, we prove Theorem 1.3 about the generic rational fibration structure of irreducible fibres on degenerations of klt Fano fibrations. First we recall the setup of this result. Fix $d\in\mathbb{N}$, $\epsilon>0$ and $t\in(0,1]$. Denote by $\mathcal{F}_{d,\epsilon,t}$ the set of data $(X/Z,tF)$ where
@@
  • $f\colon X\to Z$ is a klt Fano fibration with $\dim X=d$, that is, $X$ is a klt variety of dimension $d$ and $K_{X}$ is ample/Z/Z,
@@
Decreasing $t$ if necessary, we can assume that $t\in(0,1]$ is a rational number. Then by the existence of complements (see [Bir19, Theorem 1.8]), there exists an $n\in\mathbb{N}$ depending only on $d$ and $t$ such that there exists a monotonic $n$-complement $K_{X}+B^{+}$ of $K_{X}+tF$ over $z\in Z$. Up to shrinking $Z$ around $z$ if necessary, we can assume that the boundary $B^{+}$ is defined over the whole $Z$. More precisely, there exists a boundary $B^{+}$ on $X$ such that
@@
Note that the last condition on log discrepancy follows from that $(X,tF)$ is lc and $tF\leq B^{+}$.
@@ 4.1. Modification to generically smooth couples
There is a relatively bounded family $\mathcal{P}$ of couples $(Y/Z,D)$ such that for each $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ with an $n$-complement $K_{X}+B^{+}$ we have a couple $(Y/Z,D)\in\mathcal{P}$ admitting a birational $Z$-map $\phi\colon X\dashrightarrow Y$ such that
  • (1)
    the reduced divisor $D$ is horizontal over $Z$,
@@
Step 1. For each $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$, denote by $U_{f}$ the maximal open subset of $Z$ over which every fibre of $f$ is an $\epsilon$-lc Fano variety of dimension d1d-1. Denote by $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$ the set of all fibres of all $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. Denote by $\mathcal{F}_{d,\epsilon,t}^{\circ}$ the set of all morphisms that are the restrictions of $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. That is, every member in $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is equal to $f|_{f^{-1}(U_{f})}$ for some $f\in\mathcal{F}_{d,\epsilon,t}$. By the boundedness of $\epsilon$-lc Fano varieties of dimension d1d-1 (see [Bir21, Theorem 1.1]), we can assume that $W\to B$ is the universal family of $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$ where $W$ and $B$ are quasi-projective varieties. By taking suitable stratifications on $\mathcal{F}_{d,\epsilon,t}^{\circ}$ and doing induction on $\dim B$, we conclude that there exists a fixed natural number $\ell$ depending only on $\mathcal{F}_{d,\epsilon,t}^{\circ}$ such that for every $f\in\mathcal{F}_{d,\epsilon,t}$ there exists an open subset $U_{f}^{\prime}\subset U_{f}$ over which $\ell K_{X}$ is very ample. Without loss of generality, we can assume that $U_{f}^{\prime}$ is equal to the open subset $U_{f}$. Moreover, by the boundedness of $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$, the relative degree of $\ell K_{X}$ over $U_{f}$ is bounded from above. Thus, $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is relatively bounded.
@@
Thus, on a general fibre $X_{s}$ of $X\to Z$, we have
@@
Step 3. For every $T_{i}$, let $\widetilde{T}_{i}$ be a projective compactification of $T_{i}$. Embed $V_{i}\to T_{i}$ into some $\mathbb{P}^{n}\times T_{i}$; denote by $\widetilde{V}_{i}$ (respectively, by $\widetilde{C}_{i}$) the reduced schematic closure of $V_{i}$ (respectively, of $C_{i}$) in $\mathbb{P}^{n}\times\widetilde{T}_{i}$. Then $(\widetilde{V}_{i},\widetilde{C}_{i})\to\widetilde{T}_{i}$ is a projective compactification of $(V_{i},C_{i})\to T_{i}$ where $\widetilde{V}_{i}$, $\widetilde{C}_{i}$, and $\widetilde{T}_{i}$ are all projective. As $Z$ is a smooth curve, the morphism $U_{f}\to T_{i}$ extends to a morphism $Z\to\widetilde{T}_{i}$. Denote by $\mathcal{P}$ the set of all couples $(Y/Z,D)$ constructed from the base changes $Z\times_{\widetilde{T}_{i}}\widetilde{V}_{i}$ and $Z\times_{\widetilde{T}_{i}}\widetilde{C}_{i}$ for morphisms $Z\to\widetilde{T}_{i}$ as in Lemma 2.14. Then $\mathcal{P}$ is relatively bounded. Also, by construction, every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational $Z$-map $\phi\colon X\dashrightarrow Y$ to some $(Y/Z,D)$ in $\mathcal{P}$; moreover, $(X/Z,\operatorname{Supp}B^{+})$ and $(Y/Z,\operatorname{Supp}D)$ have the same general fibres. ∎
@@
There exists a relatively bounded family $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ of couples $(Y/Z,D)$ with $D$ horizontal/Z/Z such that
@@
    every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational $Z$-map $\phi\colon X\dashrightarrow Y$ to some $Y\to Z$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$, and
@@
      the divisor $D$ supports on the union of $\operatorname{Supp}\phi_{*}B^{+}$ and the support of all horizontal/Z/Z exceptional divisors of $\phi^{-1}$.
@@
Take the relatively bounded family of couples $\mathcal{P}$ in Lemma 4.2. Then there are finitely many universal projective morphisms of varieties as in Lemma 2.15. We can assume that there is only one such universal projective morphism $(V,C)\to T$ for $\mathcal{P}$ with $V$ and $T$ quasi-projective varieties and $C$ a reduced horizontal/T/T divisor on $V$. Take a log resolution $V^{\text{sm}}\to V$ of the couple $(V,C)$. Denote by $C^{\text{sm}}$ the reduced divisor supporting on the union of the birational transform of $C$ and all horizontal/T/T exceptional divisors of $V^{\text{sm}}\to V$.
By generic smoothness, there exists an open dense subset $U_{1}\subset T$ such that $(V^{\text{sm}},C^{\text{sm}})$ is log smooth over $U_{1}$. Moreover, we can assume that the images of all vertical/T/T exceptional divisors of $V^{\text{sm}}\to V$ do not intersect $U_{1}$. Shrinking $Z$ around the closed point $z\in Z$ if necessary, there is a morphism $Z\to T$ as in Lemma 2.15. If the image of $Z\to T$ is not entirely contained in $T_{1}\coloneqq T\setminus U_{1}$, we can take the couple $(Y/Z,D)$ as the reduction of the main component of $V^{\text{sm}}\times_{T}Z$ equipped with the reduced divisor which is the reduction of the horizontal/Z/Z part of $C^{\text{sm}}\times_{T}Z$. Write $T_{1}$ as a union of finitely many irreducible varieties, then the result follows by making a suitable stratification on $\mathcal{P}$ and doing induction on $\dim T$. ∎
@@ 4.4. Proof of bounded irrationality of irreducible fibres
Fix positive real numbers $\epsilon>0$, $t\in(0,1]$ and a natural number $d$. Assume that $f\colon X\to Z$ is a klt Fano contraction with $\dim X=d$ such that
@@
    $F$ is the reduction of an irreducible fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
@@
Decreasing $t$ if necessary, we can assume that $t$ is a rational number. Then we construct the $n$-complement $K_{X}+B^{+}$ and the family of projective couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in §4.1.
Step 1. Recall that the family of projective couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ constructed in Lemma 4.3 is relatively bounded. Denote by $(\mathfrak{X}_{i},\mathfrak{D}_{i})\to\mathfrak{B}_{i}$ the collection of universal families of couples of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in Lemma 2.15. We can assume that there is only one such projective morphism
@@
By the construction in the proof of Lemma 4.3, there exists a maximal open subset $\mathfrak{B}^{\circ}\subset\mathfrak{B}$ such that $(\mathfrak{X},\mathfrak{D})$ is log smooth over $\mathfrak{B}^{\circ}$. Moreover, after shrinking $Z$ around the closed point $z\in Z$ if necessary, we can assume that every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$ admits a morphism $Z\to\mathfrak{B}$ whose image intersects $\mathfrak{B}^{\circ}$. That is, for every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$, there is a commutative diagram for some $(Y/Z,D)\in\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as follows.
@@
If the centre of $F$ on $\mathfrak{X}$ is an irreducible component of a fibre of $\Phi$, then $F$ is birationally bounded, hence $F$ has bounded irrationality by Lemma 2.17. In this case, we take $C$ to be a resolution of the centre of $F$ on $\mathfrak{X}$. Note that given a bounded family of varieties there is a bounded family of resolutions for such varieties, hence we can take $C$ to be a bounded smooth projective variety. So, in the rest of the proof, we can assume that the centre of $F$ on $\mathfrak{X}$ is not an irreducible component of a fibre of $\Phi$.
Assume that the image of $z\in Z$ in $\mathfrak{B}$ is a closed point of $\mathfrak{B}^{\circ}$. Shrinking $Z$ around $z$ if necessary, we can assume that the whole image of $Z$ in $\mathfrak{B}$ is contained in the open subset $\mathfrak{B}^{\circ}$. Then by the construction of $\mathfrak{B}^{\circ}$, the couple $(Y,D)$ is log smooth. Write
$K_{Y}+B_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+}).$
By construction of the couple $(Y,D)$ in Lemma 4.3, we have $B_{Y}\leq D$. Thus, by Lemma 2.4,
$0\leq a(F,Y,D)\leq a(F,Y,B_{Y})=a(F,X,B^{+})<1,$
which implies that $a(F,Y,D)=0$ by Proposition 3.5; in particular, we conclude that $\operatorname{centre}_{Y}F$ is an lc centre of $(Y,D)$ contained in the fibre $Y_{z}$, so $\operatorname{centre}_{Y}F$ is a stratum of the log smooth couple $(Y_{z},D_{z})$. Then it is clear that $\operatorname{centre}_{Y}F$ is bounded. Moreover, as $(Y,D)$ is strict toroidal, we can conclude by applying Lemma 2.17 and Proposition 3.5 and by taking $C$ to be a bounded smooth resolution of $\operatorname{centre}_{Y}F$.
@@
such that $m_{\mathfrak{B}}$ and $m_{\mathfrak{X}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}}^{-1}(\mathfrak{D})$ is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$, $m_{\mathfrak{B}}^{-1}(\mathfrak{Z}_{\mathfrak{B}})$ is contained in $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$, and $\Phi^{\prime}$ is a toroidal morphism of toroidal embeddings. Moreover, without loss of generality, we can assume that the open subset $U_{\mathfrak{B}}$ in Theorem 3.12 (ii) is contained in $\mathfrak{B}^{\circ}$ of Step 1.
Let $Z\to\mathfrak{B}$ be a morphism for a couple $(Y/Z,D)$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ which corresponds to some $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$. We can assume that both $m_{\mathfrak{X}}$ and $m_{\mathfrak{B}}$ are birational over the generic point of the image of $Z$ in $\mathfrak{B}$ and assume that the image of $Z$ intersects the open subset $U_{\mathfrak{B}}\subset\mathfrak{B}$; otherwise, we make a stratification of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ and do induction on $\dim\mathfrak{B}$.
Denote by $g$ the morphism $Y\to Z$. Let $Z^{\prime}$ be the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}Z$. Then $Z^{\prime}\to Z$ is an isomorphism via $m_{\mathfrak{B}}$ as $Z$ is a smooth curve. Let $Y^{\prime}$ be the normalisation of the main component of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}Z^{\prime}$, and let $U_{Y^{\prime}}\subset Y^{\prime}$ be the pullback of $U_{\mathfrak{X}^{\prime}}$ to $Y^{\prime}$. Similarly, we can define the open subset $U_{Z^{\prime}}\subset Z^{\prime}$. Denote by $h_{Y^{\prime}}$ the induced morphism $Y^{\prime}\to\mathfrak{X}^{\prime}$ and by $g^{\prime}$ the morphism $Y^{\prime}\to Z^{\prime}$. Then by Lemma 3.7, it is clear that $U_{Y^{\prime}}=h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})$ is equal to $h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})\cap(g^{\prime})^{-1}(U_{Z^{\prime}})$. By Theorem 3.12 (ii), the induced morphism $g^{\prime}\colon(U_{Y^{\prime}}\subset Y^{\prime})\to(U_{Z^{\prime}}\subset Z^{\prime})$ is also a toroidal morphism of strict toroidal embeddings. We include the commutative diagram here for convenience.

### v3

@@ 4. Bounding irrationality of degenerations
In this section, we prove Theorem 1.3 about the generic rational fibration structure of irreducible fibres on degenerations of klt Fano fibrations. First we recall the setup of this result. Fix $d\in\mathbb{N}$, $\epsilon>0$ and $t\in(0,1]$. Denote by $\mathcal{F}_{d,\epsilon,t}$ the set of data $(X/Z,tF)$, where
@@
  • $f\colon X\to Z$ is a klt Fano fibration with $\dim X=d$, that is, $X$ is a klt variety of dimension $d$, $f$ is a contraction, and $K_{X}$ is ample/Z/Z,
@@
By decreasing $t$ if necessary, we can assume that $t\in(0,1]$ is a rational number. Then by the boundedness of complements (see [7, Theorem 1.8]), there exists an $n\in\mathbb{N}$ depending only on $d$ and $t$ such that there exists a monotonic $n$-complement $K_{X}+B^{+}$ of $K_{X}+tF$ over $z\in Z$. Up to shrinking $Z$ around $z$ if necessary, we can assume that the boundary $B^{+}$ is defined over the whole $Z$. More precisely, there exists an effective $\mathbb{Q}$-divisor $B^{+}$ on $X$ such that
@@
Note that the last condition on log discrepancy follows from $tF\leq B^{+}$. Indeed, since $F$ is a divisor on $X$, the log discrepancy $a(F,X,tF)$ is equal to 1t1-t, which is <1<1; moreover, as $tF\leq B^{+}$, we get $a(F,X,B^{+})\leq a(F,X,tF)<1$.
@@ 4.1. Modification to generically smooth couples
There is a relatively bounded family $\mathcal{P}$ of couples $(Y/Z,D)$ such that for each $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$ with an $n$-complement $K_{X}+B^{+}$ as above, we have a couple $(Y/Z,D)\in\mathcal{P}$ admitting a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ such that
@@
For each $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$, by abusing the notation slightly, we also write this data as $f\in\mathcal{F}_{d,\epsilon,t}$, where $f$ denotes the contraction $X\to Z$.
Step 1. For each $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$, denote by $U_{f}$ the maximal open subset of $Z$ over which every closed fibre of $f$ is an $\epsilon$-lc Fano variety of dimension d1d-1. Denote by $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$ the set of all closed fibres of all $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. Denote by $\mathcal{F}_{d,\epsilon,t}^{\circ}$ the set of all morphisms that are the restrictions of $f\in\mathcal{F}_{d,\epsilon,t}$ over $U_{f}$. That is, every member in $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is equal to $f|_{f^{-1}(U_{f})}$ for some $f\in\mathcal{F}_{d,\epsilon,t}$.
[…]
  • $W^{\prime}\to R^{\prime}$ is flat, where $W^{\prime}\coloneqq R^{\prime}\times_{R}W$,
  • $K_{W^{\prime}/R^{\prime}}$ is well-defined, and
  • for every closed point $p\in R^{\prime}$, if $W_{p}$ is the fibre of $W\to R$ over $p$, then $K_{W_{p}}$ is the restriction of $K_{W^{\prime}/R^{\prime}}$ to $W_{p}$.
[…]
Denote by $Z^{\prime}$ the subset of $Z$ such that $(-\ell K_{X})|_{X_{z}}$ is very ample for every $z\in Z^{\prime}$. By [24, Proposition (9.6.3)], $Z^{\prime}$ is a constructible subset of $Z$. Moreover, by assumption, $Z^{\prime}$ is an infinite subset of $Z$, then we can conclude that $Z^{\prime}$ contains a dense open subset of $Z$. In particular, $(-\ell K_{X})|_{X_{\eta_{Z}}}$ is very ample over $\operatorname{Spec}K(Z)$, where $\eta_{Z}$ is the generic point of $Z$ and $K(Z)$ is the function field of $Z$. Then $(-\ell K_{X})|_{X_{\eta_{Z}}}$ induces a closed immersion $\iota\colon X_{\eta_{Z}}\to\mathbb{P}^{N}_{K(Z)}$ for some $N\in\mathbb{N}$. By spreading out $\iota$, there is a morphism
$\iota_{Z^{\prime\prime}}\colon X\times_{Z}Z^{\prime\prime}\to\mathbb{P}^{N}_{Z^{\prime\prime}}$
over an open subset $Z^{\prime\prime}\subset U_{f}$ that extends $\iota$. Up to shrinking $Z^{\prime\prime}$, we can assume that $\iota_{Z^{\prime\prime}}$ is also a closed immersion, and that $\mathcal{O}_{X}(-\ell K_{X})$ is isomorphic to $\iota_{Z^{\prime\prime}}^{*}\mathcal{O}_{\mathbb{P}^{N}_{Z^{\prime\prime}}}(1)$ over $Z^{\prime\prime}$ by [21, 7.27, 12.93]. Thus, $\ell K_{X}$ is very ample over $Z^{\prime\prime}$. Then it suffices to take $U_{f}^{\prime}=Z^{\prime\prime}$. This concludes the claim.
Now we replace the open subset $U_{f}\subseteq Z$ by $U_{f}^{\prime}$. Then by the boundedness of $\mathcal{F}_{d,\epsilon,t}^{\text{fb}}$, the relative degree of $\ell K_{X}$ over $U_{f}$ is bounded from above. Thus, $\mathcal{F}_{d,\epsilon,t}^{\circ}$ is relatively bounded.
@@
Thus, on a general closed fibre $X_{s}$ of $X\to Z$, we have
@@
Step 3. For every $T_{i}$, let $\widetilde{T}_{i}$ be a projective compactification of $T_{i}$. Embed $V_{i}\to T_{i}$ into some $\mathbb{P}^{m}\times T_{i}$; denote by $\widetilde{V}_{i}$ (respectively, by $\widetilde{C}_{i}$) the reduced scheme-theoretic closure of $V_{i}$ (respectively, of $C_{i}$) in $\mathbb{P}^{m}\times\widetilde{T}_{i}$. Then $(\widetilde{V}_{i},\widetilde{C}_{i})\to\widetilde{T}_{i}$ is a projective compactification of $(V_{i},C_{i})\to T_{i}$, where $\widetilde{V}_{i}$, $\widetilde{C}_{i}$, and $\widetilde{T}_{i}$ are all projective. As $Z$ is a smooth curve, the morphism $U_{f}\to T_{i}$ extends to a morphism $Z\to\widetilde{T}_{i}$. Denote by $\mathcal{P}$ the set of all couples $(Y/Z,D)$ constructed from the base changes $Z\times_{\widetilde{T}_{i}}\widetilde{V}_{i}$ and $Z\times_{\widetilde{T}_{i}}\widetilde{C}_{i}$ for morphisms $Z\to\widetilde{T}_{i}$ as in Lemma 2.14, that is, $Y$ is the reduction of the main component of $Z\times_{\widetilde{T}_{i}}\widetilde{V}_{i}$, and $D$ is the reduction of the horizontal/Z/Z divisorial part of $Z\times_{\widetilde{T}_{i}}\widetilde{C}_{i}$. Then $\mathcal{P}$ is relatively bounded. Also, by construction, every $f\colon X\to Z$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ to some $(Y/Z,D)$ in $\mathcal{P}$; moreover, $(X/Z,\operatorname{Supp}B^{+})$ and $(Y/Z,\operatorname{Supp}D)$ have the same general closed fibres. ∎
@@
There exists a relatively bounded family $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ of couples $(Y/Z,D)$ such that
@@
    every $(X/Z,tF)$ in $\mathcal{F}_{d,\epsilon,t}$ admits a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ to some $(Y/Z,D)$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$, and
@@
      the divisor $D$ is supported on the union of $\operatorname{Supp}\phi_{*}B^{+}$ and the supports of all horizontal/Z/Z exceptional divisors of $\phi^{-1}$, where $K_{X}+B^{+}$ is the complement as in Lemma 4.2.
@@
Take the relatively bounded family of couples $\mathcal{P}$ in Lemma 4.2. Then there are finitely many universal projective morphisms of varieties and reduced divisors as in Lemma 2.15. We can assume that there is only one such universal projective morphism $(V,C)\to T$ for $\mathcal{P}$ with $V$ and $T$ quasi-projective varieties and $C$ a reduced horizontal/T/T divisor on $V$. Take a log resolution $V^{\text{sm}}\to V$ of the couple $(V,C)$. Denote by $C^{\text{sm}}$ the reduced divisor supported on the union of the birational transform of $C$ and all horizontal/T/T exceptional divisors of $V^{\text{sm}}\to V$.
By generic smoothness, there exists an open dense subset $U_{1}\subset T$ such that $(V^{\text{sm}},C^{\text{sm}})$ is log smooth over $U_{1}$. Moreover, we can assume that the images of all vertical/T/T exceptional divisors of $V^{\text{sm}}\to V$ do not intersect $U_{1}$. Shrinking $Z$ around the closed point $z\in Z$ if necessary, there is a morphism $Z\to T$ as in Lemma 2.15. If the image of $Z\to T$ is not entirely contained in $T_{1}\coloneqq T\setminus U_{1}$, we can take the couple $(Y/Z,D)$ as the reduction of the main component of $V^{\text{sm}}\times_{T}Z$ equipped with the reduced divisor which is the reduction of the horizontal/Z/Z divisorial part of $C^{\text{sm}}\times_{T}Z$. Write $T_{1}$ as a union of finitely many varieties. Then the result follows by induction on $\dim T$. ∎
@@ 4.4. Proof of bounded irrationality of irreducible fibres
Fix positive real numbers $\epsilon>0$, $t\in(0,1]$ and a natural number $d$. Assume that $f\colon X\to Z$ is a klt Fano fibration with $\dim X=d$ such that
@@
    $F$ is the reduction of an irreducible closed fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety depending only on $d,\epsilon,t$, hence with bounded degree of irrationality.
@@
By decreasing $t$ if necessary, we can assume that $t$ is a rational number. Then we construct the $n$-complement $K_{X}+B^{+}$ and the family of couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in §4.1, where $n\in\mathbb{N}$ depends only on $d,t$.
Step 1. Recall that the family of couples $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ constructed in Lemma 4.3 is relatively bounded. Denote by $(\mathfrak{X}_{i},\mathfrak{D}_{i})\to\mathfrak{B}_{i}$ the collection of universal families of couples of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as in Lemma 2.15. We can assume that there is only one such projective morphism
@@
By the construction in the proof of Lemma 4.3, there exists a maximal open subset $\mathfrak{B}^{\circ}\subset\mathfrak{B}$ such that $(\mathfrak{X},\mathfrak{D})$ is log smooth over $\mathfrak{B}^{\circ}$, and that every fibre of $\Phi$ over a closed point of $\mathfrak{B}^{\circ}$ is a smooth projective variety. Moreover, after shrinking $Z$ around the closed point $z\in Z$ if necessary, we can assume that every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$ admits a morphism $Z\to\mathfrak{B}$ whose image intersects $\mathfrak{B}^{\circ}$. That is, for every $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$, there is a commutative diagram for some $(Y/Z,D)\in\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ as follows,
@@
where $\phi\colon X\dashrightarrow Y$ is a birational map over $Z$. If the closure of the image of $F$ in $\mathfrak{X}$ is an irreducible component of a closed fibre of $\Phi$, then $F$ is birationally bounded, hence $F$ has bounded irrationality by Lemma 2.17. In this case, we take $C$ to be a resolution of the closure of the image of $F$ in $\mathfrak{X}$. Note that given a bounded family of varieties there is a bounded family of resolutions for such varieties, hence we can take $C$ to be a bounded smooth projective variety. So, in the rest of the proof, we can assume that the closure of the image of $F$ in $\mathfrak{X}$ is not an irreducible component of a closed fibre of $\Phi$.
Assume that the image of $z\in Z$ in $\mathfrak{B}$ is a closed point of $\mathfrak{B}^{\circ}$. Shrinking $Z$ around $z$ if necessary, we can assume that the whole image of $Z$ in $\mathfrak{B}$ is contained in the open subset $\mathfrak{B}^{\circ}$. Then by the construction of $\mathfrak{B}^{\circ}$, the couple $(Y,D)$ is log smooth. In particular, the closed fibre $Y_{z}$ of $Y\to Z$ over $z\in Z$ is a smooth projective variety, and $(Y,D+Y_{z})$ is also a log smooth couple. Write
$K_{Y}+B_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+}),$
where the coefficients of $B_{Y}$ are $\leq 1$ as $(X,B^{+})$ is lc; see Lemma 2.4. Denote by $B_{Y}^{h}$ the horizontal/Z/Z part of $B_{Y}$. By construction of the couple $(Y,D)$ in Lemma 4.3, $B_{Y}^{h}\leq D$, hence up to shrinking $Z$ around $z$, we can assume that $B_{Y}\leq D+Y_{z}$. Thus, by Lemma 2.4,
$0\leq a(F,Y,D+Y_{z})\leq a(F,Y,B_{Y})=a(F,X,B^{+})<1,$
which implies that $a(F,Y,D+Y_{z})=0$ by Proposition 3.5; in particular, we can conclude that $\operatorname{centre}_{Y}F$ is an lc centre of $(Y,D+Y_{z})$ contained in the closed fibre $Y_{z}$, so $\operatorname{centre}_{Y}F$ is a stratum of the log smooth projective couple $(Y_{z},D_{z})$. Then it is evident that $\operatorname{centre}_{Y}F$ is bounded. Moreover, as $(Y,D+Y_{z})$ is a strict toroidal couple, we can conclude by applying Lemma 2.17 and Proposition 3.5, and by taking $C$ to be a bounded smooth resolution of $\operatorname{centre}_{Y}F$.
@@
such that $m_{\mathfrak{B}}$ and $m_{\mathfrak{X}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}}^{-1}(\operatorname{Supp}\mathfrak{D})$ is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$, $m_{\mathfrak{B}}^{-1}(\mathfrak{Z}_{\mathfrak{B}})$ is contained in $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$, and $\Phi^{\prime}$ is a surjective projective toroidal morphism of strict toroidal embeddings. Moreover, without loss of generality, we can assume that the open subset $U_{\mathfrak{B}}$ in Theorem 3.12 (ii) is contained in $\mathfrak{B}^{\circ}$ of Step 1.
Let $Z\to\mathfrak{B}$ be a morphism for a couple $(Y/Z,D)$ in $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$ which corresponds to some $(X/Z,tF)\in\mathcal{F}_{d,\epsilon,t}$. We can assume that $m_{\mathfrak{B}}$ (respectively, $m_{\mathfrak{X}}$) is an isomorphism (respectively, is birational) over the generic point of the image of $Z$ in $\mathfrak{B}$ and assume that the image of $Z$ intersects the open subset $U_{\mathfrak{B}}\subset\mathfrak{B}$; otherwise, the image of $Z$ is entirely contained in a proper closed subset of $\mathfrak{B}$, then we can do induction on $\dim\mathfrak{B}$ (which corresponds to taking stratification of $\mathcal{F}_{d,\epsilon,t}^{\text{sm}}$).
Denote by $g$ the morphism $Y\to Z$. Let $Z^{\prime}$ be the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}Z$. Then $Z^{\prime}\to Z$ is an isomorphism via $m_{\mathfrak{B}}$ as $Z$ is a smooth curve. Let $Y^{\prime}$ be the normalisation of the main component of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}Z^{\prime}$, and let $U_{Y^{\prime}}\subset Y^{\prime}$ be the inverse image of $U_{\mathfrak{X}^{\prime}}$ in $Y^{\prime}$. Similarly, let $U_{Z^{\prime}}\subset Z^{\prime}$ be the inverse image of $U_{\mathfrak{B}^{\prime}}$ in $Z^{\prime}$. Denote by $h_{Y^{\prime}}$ the induced morphism $Y^{\prime}\to\mathfrak{X}^{\prime}$ and by $g^{\prime}$ the morphism $Y^{\prime}\to Z^{\prime}$. Then by Lemma 3.7, it is evident that $U_{Y^{\prime}}=h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})$ is equal to $h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})\cap(g^{\prime})^{-1}(U_{Z^{\prime}})$. By Theorem 3.12 (ii), the induced morphism $g^{\prime}\colon(U_{Y^{\prime}}\subset Y^{\prime})\to(U_{Z^{\prime}}\subset Z^{\prime})$ is also a toroidal morphism of strict toroidal embeddings. We include the commutative diagram here for convenience.